The visible light spectrum has a wavelength range of 380-700nm. Find the frequency of a wave with the wavelength of 410nm. (Hint: the speed of light is 3x10^8 m/s) 


1
Expert's answer
2021-08-16T08:32:15-0400

The frequency of a wave

"f=\frac{c}{\lambda}=\frac{3*10^8\:\rm m/s}{410*10^{-9}\:\rm m}=7.3*10^{14}\:\rm Hz"